OpenStudy (kainui):

Generalization of even/odd garbage for dan, check it @dan815

OpenStudy (shadowlegendx):

I found a garbage can

OpenStudy (kainui):

\[\LARGE \sum_{n=1}^N \sum_{k=1}^N \frac{ e^{i 2\pi n /N} f( e^{i 2\pi k /N}x)}{N}\]

OpenStudy (kainui):

f(x)=sumsum...etc...

OpenStudy (abb0t):

@sahrya

OpenStudy (kainui):

Do it for N=2 or N=1 and then see that N=3 also works as expressions for f(x).

OpenStudy (dan815):

@Heart_Broken_Kid

OpenStudy (kainui):

If you do the k summation first it will be kind of awkward. Maybe it should be flipped, so the n and k are different so that you do the other summation first.

OpenStudy (shadowlegendx):

@Obama

OpenStudy (shadowlegendx):

He will help us with this problem guys, dun worry :)

OpenStudy (anonymous):

N = 1, not equal, f(x)?

OpenStudy (kainui):

So for N=1 we have:\[\LARGE f(x)=\frac{ f(x)}{1}\] N=2 we have: \[\LARGE f(x)=\frac{f(x)+f(-x)}{2}+\frac{f(x)-f(-x)}{2}\] See how f(x) is now expressed as the sum of an even and odd function? Let's go deeper. N=3 and let's say: \[\Huge e^{i 2 \pi /3}= \omega\] \[\large f(x)=\frac{f(x)+f( \omega x)+ f(\omega^2 x)}{3}+\frac{f(x)+\omega f( \omega x)+ \omega^2 f(\omega^2 x)}{3}+ \\ \large +\frac{f(x)+\omega ^2 f( \omega x)+ \omega f(\omega^2 x)}{3}\] Make sure you can show to yourself that that's true. I hope the equation I wrote is actually describing these, they might be off slightly, but maybe you can see where I'm going with this.

OpenStudy (anonymous):

Nice nice!

OpenStudy (anonymous):

ive seen this before :o

OpenStudy (kainui):

I sort of stumbled across this when I noticed e^x is its own derivative and sinx is its own 4th derivative and wanted to find a function that was its own nth derivative. But it sort of seems like we think and talk about even and odd numbers as ways of reasoning but what about every third number reasoning? It doesn't even have a name as far as I know even/odd is 2n and 2n+1 but what about 3n, 3n+1, and 3n+2? How about even higher divisibility sort of things? Anyways just throwing around ideas haha, feel free to throw your stuff here too and talk about ideas.

OpenStudy (anonymous):

cool :D but .. according to D.A 2n , 2n+1 ( on Z) gives all sets of numbers >=2 3n,3n+1, 3n+2 gives all sets of numbers >=3 but you want know which odd which even , how ever if u used D.A for 4,6,...ect its the same to 2 but it would start from 4,6,.. so , being 2n,2n+1 do not miss any case

OpenStudy (kainui):

Well I'm not saying I'm interested in even/odd anymore. I'm saying this is like a new version of even/odd, you see?

OpenStudy (anonymous):

im interested in what u doing :o i was working on if we could see even /odd around any line hehe

OpenStudy (kainui):

what do you mean around any line? I'm curious, describe it or draw pictures.

OpenStudy (kainui):

I just noticed, so if a function is even, then it satisfies: \[\LARGE f(x)=f(-x)\] and if it's odd it satisfies: \[\LARGE f(x)=-f(-x)\] Now if we go into this new thing we have three new types of functions we can have: \[\LARGE f(x)=f( \omega x) = f( \omega^2 x)\] another is: \[\LARGE f(x)=\omega f( \omega x) = \omega ^2 f( \omega^2 x)\] and the final is \[\LARGE f(x)=\omega^2 f( \omega x) = \omega f( \omega^2 x)\] So pretty interesting, and just like all functions can be represented as the sum of an even and an odd function, so too can a function be represented as the sum of 3 functions that follow these identities.

OpenStudy (anonymous):

odd / even around any line like this